# Unveiling the Learning Mind of Language Models: A Cognitive Framework and Empirical Study

### v1

@@ 7 LearnArena: A Benchmark Suite for General Learning Ability
To assess the uniqueness and relevance of our benchmark, we compare it with a broad set of widely used evaluation benchmarks by computing Spearman rank correlations between system-level rankings. These benchmarks include instruction-following datasets such as AlpacaEval (Dubois et al., 2024), Vicuna (Chiang et al., 2023), Self-Instruct (Wang et al., 2022b), and WizardLM (Xu et al., 2023), as well as standard task-specific evaluations including ARC (Clark et al., 2018) (science reasoning), HellaSwag (Zellers et al., 2019) and Winogrande (Sakaguchi et al., 2019) (commonsense inference), GSM8K (Cobbe et al., 2021) (mathematical reasoning), and MMLU (Hendrycks et al., 2021a) (broad-domain knowledge).

### v3

@@ A.6 Details of LearnArena
To assess the uniqueness and relevance of our benchmark, we compare it with a broad set of widely used evaluation benchmarks by computing Spearman rank correlations between system-level rankings. These benchmarks include instruction-following datasets such as AlpacaEval [dubois2024alpacafarm], Vicuna [vicuna2023], Self-Instruct [wang2022self], and WizardLM [xu2023wizardlm], as well as standard task-specific evaluations including ARC [Clark2018ThinkYH] (science reasoning), HellaSwag [zellers2019hellaswag] and Winogrande [sakaguchi2019winogrande] (commonsense inference), GSM8K [cobbe2021training] (mathematical reasoning), and MMLU [hendryckstest2021] (broad-domain knowledge).